OpenStudy (anonymous):

Frank needs to convert 24 centimeters per second to meters per minute. Which conversion factors should she use?

OpenStudy (lucaz):

1m = 100cm 1min = 60s \[\huge{24\frac{cm}{s}\times \frac{1m}{100cm}\times \frac{60s}{1min}}\]